• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Solvignon

    Log entry time 08:06:05 on July23,2006

    Entry number 174960

    keyword=Shift summary

    00:05 Started shift preparing the target to run on lead.
    00:15 Controlled access to fix L-Q1. Scott did it.
    00:50 Moved target to BeO to check raster before going to lead.
    00:51 Beam is on BeO at 5uA (asked raster 5x5 to MCC, got 2x2 on the spot++.
    00:55 Asked raster 10x10.
    00:58 Beam on and we got 3x3 raster size from spot++.
    01:01 Asked raster 20x20.
    01:02 Beam on but didn't get enough stats for spot++.
    01:03 No beam.
    01:17 Beam is back. Got raster size 6x6 from spot++.
    01:20 Asked 16x16.
    01:23 Asked also 20uA in order to have radcon check the ion chamber
    radiation levels for BeO target.
    01:45 Beam is at 20uA.
    01:47 Radcon got their numbers for BeO.
    01:48 Moved target to LH2+Pb.
    01:56 Target is ready. Asked beam with raster 15x15 and current at 5uA.
    01:59 Asked 10uA.
    02:01 Asked 20uA.
    02:06 Radcon is done with the lead target test. Lead being the thickest target, other target tests are not necessary at this point.
    02:10 Move to Slanted C target for optics.
    02:18 Asked beam current at 5uA with raster off.
    02:20 Beam on.

    02:20-03:00 Run# 1563: Left at +2% and Right at 0%, slanted C --> general detector check. BPM position not so good.

    03:03 Moved target to 7 foil C.
    03:07 Hall C is ready to give us beam. So preparing for BCM calibration.

    03:10-3:30 Run# 1564: BCM calibration.

    03:31 BCM calibration is done. Asked BPM position between +/- 0.1.

    03:51-04:01 Run# 1565: Left at +2% and Right at 0%, 7 foil C --> BPM position between +/- 0.2.
    04:04-04:14 Run# 1566: Left at +2% and Right at 0%, 7 foil C --> BPM position between +/- 0.1.

    04:14 Moved left spectrometer sieve slit in.

    04:17-05:35 Run# 1567: Left at +2% and Right at 0%, 7 foil C, sieve slit.

    05:35 Moved target to slanted C.
    05:39 Asked for 20uA with raster off.

    05:42-07:00 Run# 1568: Left at +2% and Right at 0%, slanted C, sieve slit.

    07:00 Start changing Left spectrometer momentum to 0% but Gauss meter readback looks frozen. Rebooted IOC14. Moved collimator to 6msr.

    While changing left arm, taking data with right arm:
    07:03-07:19 Run# 20217: slanted C
    07:26-07:35 Run# 20218: slanted Fe (raster on, 15x15)
    07:40-07:49 Run# 20219: Ta

    07:49 Left arm is ready
    07:50 Moved target to slanted C.
    07:51 Asked 20uA raster off.
    07:54 Beam on, BPM positions look good.

    07:54-08:04 Run# 1569: Left at +0% and Right at 0%, slanted C, 6msr collimator.